# Diameter Bounds for Friends-and-Strangers Graphs

### v1

@@ Abstract.
Consider two $n$-vertex graphs $X$ and $Y$, where we interpret $X$ as a social network with edges representing friendships and $Y$ as a movement graph with edges representing adjacent positions. The friends-and-strangers graph $\mathsf{FS}(X,Y)$ is a graph on the $n!$ permutations $V(X)\to V(Y)$, where two configurations are adjacent if and only if one can be obtained from the other by swapping two friends located on adjacent positions.
Friends-and-strangers graphs were first introduced by Defant and Kravitz [6], and generalize sliding puzzles as well as token swapping problems. Previous work has largely focused on their connectivity properties. In this paper, we study the diameter of the connected components of $\mathsf{FS}(X,Y)$. Our main result shows that when the underlying friendship graph is a star with $n$ vertices, the friends-and-strangers graph has components of diameter $O(n^{4})$. This implies, in particular, that sliding puzzles are always solvable in polynomially many moves. Our work also provides explicit efficient algorithms for finding these solutions.
We then extend our results to general graphs in two ways. First, we show that the diameter is polynomially bounded when both the friendship and the movement graphs have large minimum degree. Second, when both the underlying graphs $X$ and $Y$ are Erdős-Rényi random graphs, we show that the distance between any pair of configurations is almost always polynomially bounded under certain conditions on the edge probabilities.
@@ 1. Introduction
Connectivity in $\mathsf{FS}(\mathsf{Star}_{n},Y)$ has been extensively studied starting with the work of Wilson [13] and is well understood. However, prior to our work, no general bounds were known for the diameter of connected components in the graph. The main contribution of this work is to show that sliding puzzles with $n$ tiles, whenever solvable, can be solved in $\operatorname{poly}(n)$ steps. We also show how to generalize these results to other families of friends-and-strangers graphs.
@@
Connectivity in friends-and-stangers graphs.
@@
The focus of our work is on ˜1.2, namely the diameter of friends-and-strangers graphs. Defant and Kravitz first proposed this as an open problem in their work [6, Section 7.3]. They noted that although in many settings $\mathsf{FS}(X,Y)$ is not connected, one may ask whether the diameter of each connected component in the graph is $\operatorname{poly}(n)$. (Observe that the size of $\mathsf{FS}(X,Y)$ is $n!$ – superexponential in $n$.) In other words, starting from any configuration, is it possible to get to any reachable configuration in a small number of steps?
@@ 1.1. Our Results
Our main technical result concerns $\mathsf{FS}(\mathsf{Star}_{n},Y)$, which models sliding puzzles. We show that any solvable sliding puzzle can be solved in $\operatorname{poly}(n)$ moves. In particular, the diameter of every connected component in $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is $O(n^{4})$. We also provide a family of $n$-vertex graphs $Y$ for which the diameter of $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is $\Omega(n^{3})$, showing that our upper bound is tight to within a factor of $O(n)$.
@@
Let $Y$ be a graph on $n$ vertices. Then the diameter of any connected component of $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is $O(n^{4})$.
As mentioned previously, connectivity in $\mathsf{FS}(\mathsf{Star}_{n},Y)$ was fully characterized by Wilson [13]. However, Wilson’s approach is nonconstructive and based on properties of the symmetric and alternating groups. In contrast, our results are algorithmic and constructive, providing explicit short paths between pairs of configurations in $\mathsf{FS}(\mathsf{Star}_{n},Y)$. In fact, we make strong connections between the structure of $\mathsf{FS}(\mathsf{Star}_{n},Y)$ and $\mathsf{FS}(K_{n},Y)$, showing that paths in the latter can be converted at an additional multiplicative factor of $O(n^{2})$ to paths in the former. These results are summarized in Table 1.
@@
$\mathsf{FS}(\mathsf{Star}_{n},Y)$ $O(n^{4})$ Section˜4 $\Omega(n^{3})$ Section˜4
Table 1. Diameter bounds for friends-and-strangers graphs in prior work (black) and our work (blue). These bounds refer to the largest possible diameter for any connected component of the graph.
Our next result studies the diameter of $\mathsf{FS}(X,Y)$ when both $X$ and $Y$ are Erdős-Rényi random graphs. Our investigation is inspired by the work of Alon, Defant, and Kravitz [2], who established sharp thresholds for the probability $p$ to ensure the connectivity of $\mathsf{FS}(X,Y)$ when $X$ and $Y$ are both independently drawn from $\mathcal{G}(n,p)$. We focus on a single pair of permutations over the vertex set of $X$ and determine a condition on the edge probabilities $p$ and $q$ such that the pair is connected by a short path in $\mathsf{FS}(X,Y)$ when the graphs are drawn from $\mathcal{G}(n,p)$ and $\mathcal{G}(n,q)$ respectively. This result builds upon Theorem˜1.3.
Theorem 1.4.
Let $\tau$ and $\omega$ be arbitrary permutations over $[n]$. Let $X$ and $Y$ be random graphs over the vertex set $[n]$, independently drawn from $\mathcal{G}(n,p)$ and $\mathcal{G}(n,q)$ respectively, where $p$ and $q$ satisfy $pq\geq 100\log n/n$. Then, with probability at least $1-o(n^{-2})$, the distance between $\tau$ and $\omega$ in $\mathsf{FS}(X,Y)$ is $O(n^{6})$.
Finally, we study diameter bounds for families of graphs $X,Y$ that have large minimum degree. This investigation is inspired by the work of Bangachev [3] who provided similar conditions for the connectivity of $\mathsf{FS}(X,Y)$.
Theorem 1.5.
Let $X$ and $Y$ be connected graphs on $n$ vertices such that
$\min(\delta(X),\delta(Y))+2\max(\delta(X),\delta(Y))\geq 2n.$
Then $\mathsf{FS}(X,Y)$ is connected and has diameter at most $O(n^{6})$.
Theorem 1.6.
We can visualize these results, along with the results of Bangachev [3], in Figure˜1, where we identify pairs of values of $\delta(X)$ and $\delta(Y)$ for which $\mathsf{FS}(X,Y)$ is always connected, sometimes disconnected, or always has polynomial diameter. Observe that we prove a polynomial bound on the diameter for most of the regions where Bangachev [3] proved connectivity. The only exception is the green dotted region. This leads us to conjecture that whenever $\mathsf{FS}(X,Y)$ is connected, it has polynomially bounded diameter. Resolving this conjecture is the main problem left open by our work. See Section˜7 for a discussion of open questions.
@@ 1.2. Organization of the paper
In Section˜2, we present some notation and basic claims which we will use throughout the rest of the paper. Then, in Section˜3, we bound the diameter of connected components of $\mathsf{FS}(X,Y)$ for special cases of $X$. In Section˜4, we prove our main theorem, Theorem˜1.3, and provide a lower bound on the maximum possible diameter of $\mathsf{FS}(\mathsf{Star}_{n},Y)$. In Section˜5, we prove Theorems˜1.5 and 1.6, and in Section˜6, we prove Theorem˜1.4. Finally, in Section˜7, we discuss some open directions for future work.
@@ 2.1. Notation
2.2. Preliminary observations
@@
Lemma 2.2.
Let $G$ be a biconnected graph and $e_{1}$ and $e_{2}$ be any two distinct edges in the graph. Then there is a simple (i.e., non-intersecting) cycle in $G$ containing $e_{1}$ and $e_{2}$.
Proof.
Let $e_{1}=(u_{1},v_{1})$ and $e_{2}=(u_{2},v_{2})$. Consider a graph $G^{\prime}$ obtained by replicating $G$; adding two new vertices $w_{1}$ and $w_{2}$; replacing the edge $(u_{1},v_{1})$ by the edges $(u_{1},w_{1})$ and $(w_{1},v_{1})$; and replacing the edge $(u_{2},v_{2})$ by the edges $(u_{2},w_{2})$ and $(w_{2},v_{2})$. Then, $G^{\prime}$ is also biconnected. Therefore, there exist two vertex disjoint paths from $w_{1}$ to $w_{2}$. These paths together form a cycle containing all four vertices $u_{1},u_{2},v_{1}$, and $v_{2}$. The same cycle in $G$ then contains the edges $e_{1}$ and $e_{2}$. ∎
@@ 3.1. $\mathsf{FS}(K_{n},Y)$ for general $Y$
We will first assume that $Y$ is connected. We will show that $\mathsf{FS}(K_{n},Y)$ is connected and its diameter is bounded by $\binom{n}{2}$. Let $\sigma$ be our current permutation, and $\tau$ our target permutation. For any vertex $v$ in $Y$, we’ll define $f(v)=\sigma(\tau^{-1}(v))$. This is the current location of the vertex in $K_{n}$ that we want to move to $v$. Let $S\subset V(Y)$ be initialized as the empty set. Consider the following procedure, where we will maintain the invariant that the induced subgraph on $V(Y)\setminus S$ is connected:
@@
If $Y$ is not connected, for a permutation $\sigma$ and a connected component $C$ in $Y$, let $\sigma^{-1}(C)$ denote the set of vertices of $K_{n}$ located on $C$. For two permutations $\sigma$ and $\tau$, we claim that $\sigma$ and $\tau$ are connected iff for all connected components $C$ in $Y$, $\sigma^{-1}(C)=\tau^{-1}(C)$. On the one hand, if the latter condition does not hold, then there exists a vertex $v$ in $K_{n}$ that is located at different components in $Y$ under $\sigma$ and $\tau$. But no friendly move can move the vertex from one component to another, so $\sigma$ and $\tau$ cannot be connected. On the other hand, if the condition holds, then we can transform $\sigma$ to $\tau$ by transforming the permutation over each connected component in sequence. The number of steps taken, when $Y$ has connected components $C_{1},\ldots,C_{k}$, is $\sum_{i\leq k}{{|C_{i}|}\choose 2}\leq{n\choose 2}$.
We will now show that equality holds in Theorem˜3.1 when $Y$ is isomorphic to $\mathsf{Path}_{n}$. We’ll number the vertices in $K_{n}$ $1,\cdots,n$, and number the vertices in $\mathsf{Path}_{n}$ $1,\cdots n$. Let $\sigma$ be the permutation $n,n-1,\cdots,1$ of the vertices of $K_{n}$ onto $\mathsf{Path}_{n}$, and let $\tau$ be the permutation $1,2,\ldots,n$. Define an inversion in a permutation $\omega$ to be a pair $(i,j)$ of vertices in $K_{n}$ such that $i>j$ but $\omega(i)<\omega(j)$.

### v2

@@ Abstract.
Consider two $n$-vertex graphs $X$ and $Y$, where we interpret $X$ as a social network with edges representing friendships and $Y$ as a movement graph with edges representing adjacent positions. The friends-and-strangers graph $\mathsf{FS}(X,Y)$ is a graph on the $n!$ permutations $V(X)\to V(Y)$, where two configurations are adjacent if and only if one can be obtained from the other by swapping two friends located on adjacent positions. Friends-and-strangers graphs were first introduced by Defant and Kravitz, and generalize sliding puzzles as well as token swapping problems. Previous work has largely focused on their connectivity properties.
In this paper, we study the diameter of the connected components of $\mathsf{FS}(X,Y)$. We extend the result of Kornhauser, Miller, and Spirakis on sliding puzzles to general graphs in two ways. First, we show that the diameter of $\mathsf{FS}(X,Y)$ is polynomially bounded when both the friendship and the movement graphs have large minimum degree. Second, when both the underlying graphs $X$ and $Y$ are Erdős-Rényi random graphs, we show that the distance between any pair of configurations is almost always polynomially bounded under certain conditions on the edge probabilities.
@@ 1. Introduction
Connectivity in $\mathsf{FS}(\mathsf{Star}_{n},Y)$ has been extensively studied starting with the work of Wilson [13] and is well understood. Additionally, general bounds for the diameter of connected components in this graph were proven by Kornhauser, Miller, and Spirakis [7]. The main contribution of this work is to use the latter result to prove diameter bounds for other families of friends-and-strangers graphs.
@@
Connectivity in friends-and-strangers graphs.
@@
The focus of our work is on ˜1.2, namely the diameter of friends-and-strangers graphs. Defant and Kravitz first proposed this as an open problem in their work [5, Section 7.3]. They noted that although in many settings $\mathsf{FS}(X,Y)$ is not connected, one may ask whether the diameter of each connected component in the graph is $\operatorname{poly}(n)$. (Observe that the size of $\mathsf{FS}(X,Y)$ is $n!$ – superexponential in $n$.) In other words, starting from any configuration, is it possible to get to any reachable configuration in a small number of steps?
@@ 1.1. Our Results
Our main contribution is establishing diameter bounds for families of graphs $X,Y$ that have large minimum degree. This investigation is inspired by the work of Bangachev [2] who provided similar conditions for the connectivity of $\mathsf{FS}(X,Y)$.
@@
Let $X$ and $Y$ be connected graphs on $n$ vertices such that
$\min(\delta(X),\delta(Y))+2\max(\delta(X),\delta(Y))\geq 2n.$
Then $\mathsf{FS}(X,Y)$ is connected and has diameter at most $O(n^{5})$.
Theorem 1.4.
We can visualize these results, along with the results of Bangachev [2], in Figure˜1, where we identify pairs of values of $\delta(X)$ and $\delta(Y)$ for which $\mathsf{FS}(X,Y)$ is always connected, sometimes disconnected, or always has polynomial diameter. Observe that we prove a polynomial bound on the diameter for most of the regions where Bangachev [2] proved connectivity. The only exception is the green dotted region. This leads us to conjecture that whenever $\mathsf{FS}(X,Y)$ is connected, it has polynomially bounded diameter. Resolving this conjecture is the main problem left open by our work. See Section˜6 for a discussion of open questions.
Our next result studies the diameter of $\mathsf{FS}(X,Y)$ when both $X$ and $Y$ are Erdős-Rényi random graphs. Our investigation is inspired by the work of Alon, Defant, and Kravitz [1], who established sharp thresholds for the probability $p$ to ensure the connectivity of $\mathsf{FS}(X,Y)$ when $X$ and $Y$ are both independently drawn from $\mathcal{G}(n,p)$. We focus on a single pair of permutations over the vertex set of $X$ and determine a condition on the edge probabilities $p$ and $q$ such that the pair is connected by a short path in $\mathsf{FS}(X,Y)$ when the graphs are drawn from $\mathcal{G}(n,p)$ and $\mathcal{G}(n,q)$ respectively. This result builds upon the $\mathsf{FS}(\mathsf{Star}_{n},Y)$ result in [7].
Theorem 1.5.
Let $\tau$ and $\omega$ be arbitrary permutations over $[n]$. Let $X$ and $Y$ be random graphs over the vertex set $[n]$, independently drawn from $\mathcal{G}(n,p)$ and $\mathcal{G}(n,q)$ respectively, where $p$ and $q$ satisfy $pq\geq 100\log n/n$. Then, with probability at least $1-o(n^{-2})$, the distance between $\tau$ and $\omega$ in $\mathsf{FS}(X,Y)$ is $O(n^{5})$.
Several of our results are shown in Table˜1.
@@
$\mathsf{FS}(X,Y)$ where $x>y$
and $2x+y>2n$ $O(n^{5})$ Unknown
$\mathsf{FS}(X,Y)$ where $x>y$
and $x+y>3n/2$ $O(n^{2})$ Unknown
Table 1. Diameter bounds for friends-and-strangers graphs in prior work (black) and our work (blue). These bounds refer to the largest possible diameter for any connected component of the graph. We refer to the minimum degree of $X$ as $x$ and the minimum degree of $Y$ as $y$.
@@ 1.2. Organization of the paper
In Section˜2, we present some notation and basic claims which we will use throughout the rest of the paper. Then, in Section˜3, we bound the diameter of connected components of $\mathsf{FS}(X,Y)$ for special cases of $X$. In Section˜4, we prove Theorems˜1.3 and 1.4, and in Section˜5, we prove Theorem˜1.5. Finally, in Section˜6, we discuss some open directions for future work.
@@ 2.1. Notation
2.2. Preliminary observations and past results
@@
The following result is the aforementioned $\mathsf{FS}(\mathsf{Star}_{n},Y)$ result from [7].
Theorem 2.2.
Let $Y$ be a graph on $n$ vertices. Then, the diameter of any connected component of $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is $O(n^{3})$.
This result forms the basis of several of our results later in this paper. We next present Wilson’s result on the connectivity of $\mathsf{FS}(\mathsf{Star}_{n},Y)$.
Theorem 2.3 ([13]).
@@ 3.1. $\mathsf{FS}(K_{n},Y)$ for general $Y$
We will first assume that $Y$ is connected. We will show that $\mathsf{FS}(K_{n},Y)$ is connected and its diameter is bounded by $\binom{n}{2}$. Let $\sigma$ be our current permutation, and $\tau$ our target permutation. For any vertex $v$ in $Y$, we will define $f(v)=\sigma(\tau^{-1}(v))$. This is the current location of the vertex in $K_{n}$ that we want to move to $v$. Let $S\subset V(Y)$ be initialized as the empty set. Consider the following procedure, where we will maintain the invariant that the induced subgraph on $V(Y)\setminus S$ is connected:
@@
If $Y$ is not connected, for a permutation $\sigma$ and a connected component $C$ in $Y$, let $\sigma^{-1}(C)$ denote the set of vertices of $K_{n}$ located on $C$. For two permutations $\sigma$ and $\tau$, we claim that $\sigma$ and $\tau$ are connected if and only if for all connected components $C$ in $Y$, $\sigma^{-1}(C)=\tau^{-1}(C)$. On the one hand, if the latter condition does not hold, then there exists a vertex $v$ in $K_{n}$ that is located at different components in $Y$ under $\sigma$ and $\tau$. But no friendly move can move the vertex from one component to another, so $\sigma$ and $\tau$ cannot be connected. On the other hand, if the condition holds, then we can transform $\sigma$ to $\tau$ by transforming the permutation over each connected component in sequence. The number of steps taken, when $Y$ has connected components $C_{1},\ldots,C_{k}$, is $\sum_{i\leq k}\binom{\left\lvert C_{i}\right\rvert}{2}\leq\binom{n}{2}$.
We will now show that equality holds in Theorem˜3.1 when $Y$ is isomorphic to $\mathsf{Path}_{n}$. We will number the vertices in $K_{n}$ $1,\cdots,n$, and number the vertices in $\mathsf{Path}_{n}$ $1,\cdots n$. Let $\sigma$ be the permutation $n,n-1,\cdots,1$ of the vertices of $K_{n}$ onto $\mathsf{Path}_{n}$, and let $\tau$ be the permutation $1,2,\ldots,n$. Define an inversion in a permutation $\omega$ to be a pair $(i,j)$ of vertices in $K_{n}$ such that $i>j$ but $\omega(i)<\omega(j)$.